# All-Optical Brillouin Random number Generator

### v2

@@ Acknowledgments
VII Appendix
To pass all NIST tests we generate a sample of 396.308 bits using numerical simulation of stochastic Eqns. (2)-(4). The simulation parameters are the same as in Fig. 2 except the pump amplitude $\Omega_{1}=7.8294\cdot\Omega_{th}$. Each NIST statistical test characterized by the P-value which correspond to the probability that a RNG would have produced a sample less random than the sample that was tested [NIST]. If the P-value is equal zero than the sample is completely non-random, if the P-value is equal one than the sample is absolutely random. To complete the test we use NISTs’ recommendation to take a significance level $\alpha=0.01$, when P-value >0.01>0.01 the test is passed. Some tests required a minimal bit sample to pass a test (e.g., Mauerer’s Universal Test [NIST]), which required generating the corresponding number of bits. Fig.  8 demonstrates tests results for the sample of 396.308 bits, for each test the P-value is greater than 0.01, it means that the bit sample passed all 15 NIST statistical test.

### v3

@@ Acknowledgments
Appendix A Stochastic equation simulation method
In our numerical simulation we model the following system of stochastic equation
$d{a_{1}}=\left(\left({-{\gamma_{1}}-i{\delta\omega_{1}}}\right){a_{1}}-ig{a_{2}}b-i\Omega_{1}\right)dt$ (11)
${d{a_{2}}}=\left(-\gamma_{2}{a_{2}}-i{g}{a_{1}}{b^{*}}-i\Omega_{2}\right)dt$ (12)
$\begin{gathered}{db}=\left(\left({-{\gamma_{b}}-i(\omega_{b}+\delta\omega_{2})}\right)b-i{g}{a_{1}}a_{2}^{*}\right)dt\\ +\sqrt{\gamma_{b}\bar{n}}\left(dW_{1}+idW_{2}\right)\end{gathered}$ (13)
The thermal noise $\xi(t)$ is implemented as the formal derivative of a Wiener process $\xi(t)=\xi_{1}(t)+i\xi_{2}(t)=(dW_{1}+idW_{2})/dt$ with correlation
$\left\langle dW_{1,2}(t)dW_{1,2}(t^{\prime})\right\rangle=\begin{cases}\sigma^{2}dt,&t=t^{\prime}\\ 0,&t\neq t^{\prime}\end{cases}$
[…]
Appendix B States lifetime distribution and error
Figure 8: The distribution of lifetimes for generating state. The pump amplitude is $\Omega_{1}=7.9\cdot 10^{-1}\Omega_{th}$.
In our simulations, the state lifetimes $\tau_{g/ng}$ is defined as the average value over a set of a generated sample
$\tau_{g,ng}=\frac{1}{N}\sum_{n=1}^{N}T^{n}_{g,ng}$
Here $T^{n}_{g,ng}$ is the $n$-th value of the lifetime of the generating and the non-generating states. We note that the lifetimes follow an exponential distribution (Fig. 8), with the mean of the lifetime distribution equal to its square root of variance. The standard error for such distribution can be estimated as $\Delta\tau_{g/ng}=\tau_{g,ng}/\sqrt{N}$.
Figure 9: The dependence of the standard error of the average lifetimes of the non-generating state, $\Delta\tau_{ng}$, (a) and the generating, $\Delta\tau_{g}$, (b) on the pump amplitude $\Omega_{1}$ and $\Omega_{2}$. The parameters are the same as in Fig. 2.
From the Fig. 9 the error for the non-generating state is in the range $0<\Delta\tau_{ng}<4.9/\gamma_{2}$ and the generating state is in the range $0<\Delta\tau_{g}<12.3/\gamma_{2}$.
Appendix C NIST Verification Test
To pass all NIST tests we generate a sample of 396.308 bits using numerical simulation of stochastic Eqns. (2)-(4). The simulation parameters are the same as in Fig. 2 except the pump amplitude $\Omega_{1}=7.8294\cdot\Omega_{th}$. Each NIST statistical test characterized by the P-value which correspond to the probability that a RNG would have produced a sample less random than the sample that was tested [24]. If the P-value is equal zero than the sample is completely non-random, if the P-value is equal one than the sample is absolutely random. To complete the test we use NISTs’ recommendation to take a significance level $\alpha=0.01$, when P-value >0.01>0.01 the test is passed. Some tests required a minimal bit sample to pass a test (e.g., Mauerer’s Universal Test [24]), which required generating the corresponding number of bits. Fig.  10 demonstrates tests results for the sample of 396.308 bits, for each test the P-value is greater than 0.01, it means that the bit sample passed all 15 NIST statistical test.